A home seller was offered $320,000: $140,000 cash plus zero-coupon bonds paying $180,000 after 20 years. If the discount rate is 10%, what is the buyer’s actual cost of the property?
Answer #1

Actual cost = initial cost+present value of the zero coupon bonds

Present value = 180000/(1+0.10)^20 = 180000/7.727499 = 23293.4356

Actual cost of the property = 140000+23293.4356 = 163293.44